OpenStudy (anonymous):

f(x) = (x+3) (x-4) + 4 when divided by (x-k) the remainder is k. find k?.. how do i find k..?

OpenStudy (anonymous):

the 1st step is to chos value of k from given f(x) i.e. either k=-3 or k=4 put these values simultaneously at the place of k and divide it...say if k=4 then.. put it in (x-k)=(x-4) then (x+3)times (x-4)=(x+3)(x-4) now subtract it from f(x),you will get the reminder..i.e. 4 so k=4.. ..!!

OpenStudy (anonymous):